OpenStudy (anonymous):

[8.03] Determine the factors of x2 − 14x + 24. (x + 12)(x + 2) (x + 3)(x + 8) (x − 12)(x − 2) (x − 3)(x − 8)

OpenStudy (anonymous):

(x-2)(x-12)

OpenStudy (anonymous):

how did you get that can you me step by step

OpenStudy (luigi0210):

Stop giving away answers

OpenStudy (anonymous):

what two numbers can u add together to give you -14 but when you multiply those two numbers give you 24?

OpenStudy (anonymous):

it's -12 and -2

OpenStudy (anonymous):

and that's how you get (x − 12)(x − 2)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

-2-12=-14 -2*-12=24

OpenStudy (anonymous):

you can apply this to almost any quadratic

OpenStudy (anonymous):

give me thumbs up if you like my post!